OpenStudy (anonymous):

PLEASE PLEASE HELP THANK YOU If P(x1) and P(x2) are points in quadrant II on the terminal side of the angles x1 and x2, respectively, with cos(x1) = − 1/5 and sin(x2) = 2/5, find sin(x1 + x2), cos(x1 + x2), sin(x1 − x2), and cos(x1 − x2).

OpenStudy (xapproachesinfinity):

LET'S TRY CONSTRUCTING THAT DIAGRAM FIRST

OpenStudy (xapproachesinfinity):

 WHICH IS QUADRANT II

OpenStudy (anonymous):

the top left

OpenStudy (xapproachesinfinity):

HMM LET ME THINK ABOUT THIS TERMINAL SIDE THING WHAT IT MEANS

OpenStudy (xapproachesinfinity):

EH FORGET THE DIAGRAM WE CAN DO THIS DIRECTLY WE JUST NEED TO KNOW THAT \[\sin (a\pm b)=\sin a \cos b\pm\sin b \cos a\] and \[\cos (a+b)=\cos(a)\cos(b)-\sin (a) \cos(b)\\ \cos (a-b)=\cos(a)\cos(b)+\sin (a) \cos(b)\]

OpenStudy (xapproachesinfinity):

this is all you need to do this a=x1 b=x2 they give you the info you need

OpenStudy (anonymous):

but theyre not on the unit circle. I dont understand that part

OpenStudy (xapproachesinfinity):

what is not on the unit circle?

OpenStudy (anonymous):

like for instance sin (-1/5)

OpenStudy (xapproachesinfinity):

what are talking about we have sin(x1)=-1/5 you don't need sin(-1/5)

OpenStudy (anonymous):

yeah but I just dont get how to do each opration...

OpenStudy (anonymous):

like for sin(x1+x2)...I dont get what I do to get that answer....

OpenStudy (xapproachesinfinity):

i just give the way i used a and b instead of x1 and x2

OpenStudy (anonymous):

but what do you put in for a and b?

OpenStudy (xapproachesinfinity):

you put x1 and x2 you don't need to put values

OpenStudy (xapproachesinfinity):

the values you gonna need to put are for sin(x1) sin(x2) , cos(x1), cos(x2)

OpenStudy (xapproachesinfinity):

but i just realized they didn't gave you sin(x1) and cos(x2) they only gave you cos(x1) and sin(x2) which brings us to what started with we actually need to draw that figure to find them

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

this should be something that the figure should look like so by pythagorean we have the opposite side to x1 = root(25-1)=root(24) adjacent side to x2=root(25-4)=root(21) so sin(x1)= root(24)/5 cos(x2)=root(21)/5 now it is complete